# ChatGPT’s One-year Anniversary: Are Open-Source Large Language Models Catching up?

### 28 Nov 2023 v1

@@ 3.2 Agent Capabilities
With the recent advancements in scaling up model size, LLM-based agents have drawn a great deal of attention from the NLP community. In light of this, we investigate the agent capabilities of open-source LLMs on a variety of benchmarks. Depending on the skills required, existing benchmarks can be mainly divided into four categories.
  • Using Tools: Some benchmarks have been proposed to evaluate the tool usage capabilities of LLMs. API-Bank (Li et al., 2023b) is specifically designed for tool-augmented LLMs. ToolBench (Xu et al., 2023c) is a tool manipulation benchmark including various software tools for real-world tasks. APIBench (Patil et al., 2023) consists of APIs from HuggingFace, TorchHub, and TensorHub. ToolAlpaca (Tang et al., 2023a) develops a diverse and comprehensive tool-use dataset through a multi-agent simulation environment. Coincidentally, another instruction-tuning dataset constructed using ChatGPT for tool use is also named ToolBench (Qin et al., 2023b). Besides, MINT (Wang et al., 2023d) can evaluate the proficiency of LLMs in employing tools to solve tasks that necessitate multi-turn interactions.
  • Self-Debugging: Several datasets are available to assess the ability of LLMs to self-debug, including InterCode-Bash and InterCode-SQL (Yang et al., 2023b), MINT-MBPP and MINT-HumanEval (Wang et al., 2023d), and RoboCodeGen (Liang et al., 2023).
  • Following Natural Language Feedback: MINT (Wang et al., 2023d) can also be used to measure the ability of LLMs to leverage natural language feedback by using GPT-4 (OpenAI, 2023b) to simulate human users.
  • Exploring Environment: ALFWorld (Shridhar et al., 2020), InterCode-CTF (Yang et al., 2023b), and WebArena (Zhou et al., 2023c) are introduced to evaluate whether LLMs-based agents are able to gather information from the environment and make decisions.
@@
ALFWorld IC-CTF WebAreana Code Generation
@@
Models
By pre-training Llama-2 using a code-intensive corpus containing 90B tokens and instruction fine-tuning on 300K examples including both text and code, Lemur-70B-chat (Xu et al., 2023d) surpasses the performance of GPT-3.5-turbo when exploring the environment or following natural language feedback on coding tasks. AgentTuning (Zeng et al., 2023) conducts instruction tuning with Llama-2 on the combination of its constructed AgentInstruct dataset and general domain instructions, resulting in AgentLlama. Notably, AgentLlama-70B achieves comparable performance to GPT-3.5-turbo on unseen agent tasks. Through fine-tuning Llama-2-7B on ToolBench, ToolLLaMA (Qin et al., 2023b) demonstrates comparable performance to GPT-3.5-turbo in tool usage evaluations. Chen et al. (2023a) introduce FireAct, which can fine-tune Llama-2-13B to outperform prompting GPT-3.5-turbo on HotpotQA (Yang et al., 2018). In addition, Gorilla (Patil et al., 2023), fine-tuned from Llama-7B, outperforms GPT-4 on writing API calls.
3.3 Logical Reasoning Capabilities
@@
Logical reasoning serves as fundamental capability of high-level ability and skill, such as programming, theorem proving, as well as arithmetic reasoning. To this end, in this section, we will cover the following benchmarks:
@@
  • TheoremQA (Wenhu et al., 2023) is a theorem-driven question answering dataset designed to evaluate AI models’ capabilities to apply theorems to solve challenging science problems. TheoremQA is curated by domain experts containing 800 high-quality questions covering 350 theorems from Math, Physics, EE&CS, and Finance.
  • HumanEval (Chen et al., 2021) is a set of 164 hand written programming problems. Each problem includes a function signature, docstring, body, and several unit tests, with an average of 7.7 tests per problem.
@@
  • APPs (Hendrycks et al., 2021a) is a benchmark for code generation measuring the ability of models to take an arbitrary natural language specification and generate satisfactory Python code. The benchmark includes 10,000 problems, which range from having simple one line solutions to being substantial algorithmic challenges.
@@
Different from conventional knowledge distillation based instruction tuning, Luo et al. (2023c, a) employed Evol-Instruct (Xu et al., 2023a) to construct the task-specific high quality instruction tuning dataset, where the seed instructions have evolved to the ones either extended in knowledge boundary or the depth of task complexity. Besides, Luo et al. (2023a) also incorporate PPO (Schulman et al., 2017a) algorithm to further improve the quality of both generated instruction and answer. After obtaining the expanded instruction pool, the new instruction tuning dataset is generated by collecting responses from another LLM, e.g., GPT-3.5-turbo. Finally, benefiting from the evolved depth and width of queries, the fine-tuned model achieves even better performance than GPT-3.5-turbo. For example, WizardCoder (Luo et al., 2023c) outperforms GPT-3.5-turbo on HumanEval with 19.1% absolute improvements. And WizardMath (Luo et al., 2023a) has also obtained 42.9% absolute improvements compared with GPT-3.5-turbo.
@@
Lemur (Xu et al., 2023d) has verified a better mixture of natural language data and code and induces the LLMs with stronger abilities on function calling, automatic programming, and agent. Specifically, Lemur-70B achieves significant improvements over GPT-3.5-turbo on both HumanEval and GSM8K without task-specific fine-tuning. Phi (Gunasekar et al., 2023; Li et al., 2023e) takes a different road by using the textbook as the main corpus for pre-training, which makes the strong abilities observable on much smaller language models.
@@ 3.4 Modelling Long-context Capabilities
  • SCROLLS (Shaham et al., 2022) is a popular evaluation benchmark made of 7 datasets with naturally long input. The tasks cover summarization, question-answering and natural language inference.
  • ZeroSCROLLS (Shaham et al., 2023) builds on SCROLLS (discarding ContractNLI, reusing the 6 other datasets, and adding 4 datasets) and only considers the zero-shot setting, evaluating LLMs out-of-the-shelf.
  • LongBench (Bai et al., 2023) sets a bilingual English/Chinese long-context benchmark of 21 datasets across 6 tasks.
@@
Models
On LongBench, L-Eval, BAMBOO and M4LE benchmarks, GPT-3.5-turbo or its 16k version largely outperform all open-source LLMs, such as Llama-2, LongChat, or Vicuna ; showing that it is not trivial to drive up the performance of open-source LLM on long-input tasks. Llama-2-long (Xiong et al., 2023) continues pre-training of Llama-2 with 400B tokens using a 16k context window (up from 4k in Llama-2). The resulting Llama-2-long-chat-70B outperforms GPT-3.5-turbo-16k by 37.7 to 36.7 on ZeroSCROLLS. Approaches to tackle long-context tasks include context window extension with positional interpolation (Chen et al., 2023d), which involves another (short) round of fine-tuning with longer context window ; and retrieval augmentation (Lewis et al., 2020), which necessitates access to a retriever to find relevant information. (Xu et al., 2023b) combines both these seemingly opposite techniques, pushing a Llama-2-70B above GPT-3.5-turbo-16k on average over 7 long-context tasks (including 4 datasets from ZeroSCROLLS).
@@ 3.5.1 Query-focused Summarization
Query-focused or aspect-based summarization requires to generate summaries with regard to a fine-grained question or an aspect category. Query-focused datasets include AQualMuse (Kulkarni et al., 2020), QMSum (Zhong et al., 2021) and SQuALITY (Wang et al., 2022), while Aspect-based datasets include CovidET (Zhang et al., 2023a), NEWTS (Bahrainian et al., 2022), WikiAsp (Hayashi et al., 2021), etc.
Models
(Yang et al., 2023d) finds that standard fine-tuning on training data is still better in performance compared to ChatGPT, with an average of 2 points ROUGE-1 improvement over CovidET, NEWTS, QMSum and SQuALITY.
@@ 3.5.2 Open-ended QA
Models
InstructRetro (Wang et al., 2023a) shows large improvement over GPT-3 on NQ, TriviaQA, SQuAD 2.0 and DROP, while having 7-10 percent improvement compared to a proprietary GPT-instruct model of similar size, over a range of short-form and long-form open-ended QA datasets. Initialized from a pretrained GPT model, InstructRetro continues pretraining with retrieval and then undergoes instruction tuning. [^7] InstructRetro is not yet open-sourced.
@@ 3.5.3 Medical
For mental health, IMHI (Yang et al., 2023c) benchmark is constructed using 10 existing mental health analysis datasets, including mental health detection: DR (Pirina & Çöltekin, 2018), CLP (Coppersmith et al., 2015), Dreaddit (Turcan & McKeown, 2019), loneliness, SWMH and T-SID x (Ji et al., 2022); mental health cause detection: SAD (Mauriello et al., 2021), CAMS (Garg et al., 2022); mental risk factors detection: MultiWD (SATHVIK & Garg, 2023), IRF (Garg et al., 2023).
@@
Models
For mental health, MentalLlama-chat-13B (Yang et al., 2023c) finetunes a Llama-chat-13B model on IMHI training set. MentalLlama-chat-13B model with zero-shot prompting outperforms ChatGPT with few-shot prompting or with zero-shot prompting for 9 out of 10 tasks in IMHI. Liu et al. (2023) proposes to fine-tune a Llama checkpoint to generate impression text given radiology report findings. The resulting Radiology-Llama-2 model outperforms ChatGPT and GPT-4 by a large margin on both MIMIC-CXR and OpenI datasets.
@@ 3.5.4 Generating Structured Responses
Rotowire (Wiseman et al., 2017) contains NBA game summaries with corresponding score tables. Struc-Bench (Tang et al., 2023b) introduces two datasets: Struc-Bench-Latex of which the outputs are tables in Latex format, and Struc-Bench-HTML with outputs as tables in HTML format.
Models
@@ 3.5.5 Generating Critiques
Shepherd (Wang et al., 2023c) is a 7B model initialized from Llama-7B and trained on community collected critique data and 1,317 examples of high quality human annotated data. Shepherd generates critiques on a range of diverse NLP datasets: AlpacaFarm, FairEval, CosmosQA (Huang et al., 2019), OBQA (Mihaylov et al., 2018a), PIQA (Bisk et al., 2020), TruthfulQA and CritiqueEval. With GPT-4 as an evaluator, Shepherd wins or equals ChatGPT over 60% of the time. With human evaluators, Shepherd is almost on-par with ChatGPT.
@@ 3.6 Towards Trust-worthy AI
To ensure LLMs can be trusted by humans in real-world applications, an important consideration is their reliability. For example, concerns on hallucination (Ye & Durrett, 2022; Zhao et al., 2023) and safety (Zhiheng et al., 2023b) could deteriorate user trust in LLMs and lead to risks in high-impact applications.
@@ 3.6.1 Hallucination
  • HaluEval (Li et al., 2023a) is a large dataset of generated and human-annotated hallucinated samples. It spans three tasks: question answering, knowledge-grounded dialogue, and text summarization.
@@
There exist several existing surveys on hallucination (Zhang et al., 2023b; Rawte et al., 2023) that investigate potential methodologies in detail. Specifically, the methods that surpass the current GPT-3.5-turbo performance can be either incorporated during fine-tuning or only at inference time. Selected performance metrics are shown in Table 3.
@@
During inference, existing techniques include specific decoding strategies, external knowledge augmentation, and multi-agent dialogue. For decoding, Dhuliawala et al. (2023) introduces Chain-of-Verification (CoVe), where the LLM drafts verification questions and self-verify the responses. CoVe leads to a substantial improvement on FactScore over GPT-3.5-turbo.
For external knowledge augmention, various frameworks incorporate different searching and prompting techniques to the current improve GPT-3.5-turbo performance. Li et al. (2023c) designs Chain-of-Knowledge (CoK), which retrieves from heterogenous knowledge sources before answering. Peng et al. (2023) proposes LLM-AUGMENTER, which augments LLMs with a set of plug-and-play modules and iteratively revises LLM prompts to improve model responses using feedback generated by utility functions. Knowledge Solver (KSL) (Feng et al., 2023) tries to teach LLMs to search for essential knowledge from external knowledge bases by harnessing their own strong generalizability. CRITIC (Gou et al., 2023) allows LLM to validate and progressively amend their own outputs in a manner similar to human interaction with tools. Luo et al. (2023b) introduces a Parametric Knowledge Guiding (PKG) framework, which equips LLMs with a knowledge-guiding module to access relevant knowledge without altering the LLMs’ parameters. These inference techniques then improve answer accuracy compared to the naive prompting strategy using GPT-3.5-turbo. Currently, GPT-3.5-turbo has also incorporated a retrieval plugin (OpenAI, 2023a) that accesses external knowledge to reduce hallucinations.
For multi-agent dialogue, Cohen et al. (2023) facilitates a multi-turn interaction between the Examinee LLM that generated the claim and another Examiner LLM which introduces questions to discover inconsistencies. Through the cross-examination process, performance on various QA tasks is improved. Du et al. (2023) asks multiple language model instances to propose and debate their individual responses and reasoning processes over multiple rounds to arrive at a common final answer, which improves on multiple benchmarks.

### 15 Jan 2024 v4

@@ 3.2 Agent Capabilities
With the recent advancements in scaling up model size, LLM-based agents (also called language agents) have drawn a great deal of attention from the AI community. In light of this, we investigate the agent capabilities of open-source LLMs on a variety of benchmarks. Depending on the skills required, existing benchmarks can be mainly divided into four categories.
  • Using Tools: Some benchmarks have been proposed to evaluate the tool usage capabilities of LLMs. API-Bank (Li et al., 2023c) is specifically designed for tool-augmented LLMs. ToolBench (Xu et al., 2023c) is a tool manipulation benchmark including various software tools for real-world tasks. APIBench (Patil et al., 2023) consists of APIs from HuggingFace, TorchHub, and TensorHub. ToolAlpaca (Tang et al., 2023a) develops a diverse and comprehensive tool-use dataset through a multi-agent simulation environment. Coincidentally, another instruction-tuning dataset constructed using ChatGPT for tool use is also named ToolBench (Qin et al., 2023b). Besides, MINT (Wang et al., 2023f) can evaluate the proficiency of LLMs in employing tools to solve tasks that necessitate multi-turn interactions.
  • Self-Debugging: Several datasets are available to assess the ability of LLMs to self-debug, including InterCode-Bash and InterCode-SQL (Yang et al., 2023c), MINT-MBPP and MINT-HumanEval (Wang et al., 2023f), and RoboCodeGen (Liang et al., 2023).
  • Following Natural Language Feedback: MINT (Wang et al., 2023f) can also be used to measure the ability of LLMs to leverage natural language feedback by using GPT-4 (OpenAI, 2023b) to simulate human users.
  • Exploring Environment: ALFWorld (Shridhar et al., 2020), InterCode-CTF (Yang et al., 2023c), and WebArena (Zhou et al., 2023b) are introduced to evaluate whether LLMs-based agents are able to gather information from the environment and make decisions.
Performance of LLMs
By pre-training Llama-2 using a code-intensive corpus containing 90B tokens and instruction fine-tuning on 300K examples including both text and code, Lemur-70B-chat (Xu et al., 2023d) surpasses the performance of GPT-3.5-turbo when exploring the environment or following natural language feedback on coding tasks.
@@
ALFWorld IC-CTF WebArena Code Generation
@@
AgentTuning (Zeng et al., 2023) conducts instruction tuning with Llama-2 on the combination of its constructed AgentInstruct dataset and general domain instructions, resulting in AgentLlama. Notably, AgentLlama-70B achieves comparable performance to GPT-3.5-turbo on unseen agent tasks. Through fine-tuning Llama-2-7B on ToolBench, ToolLLaMA (Qin et al., 2023b) demonstrates comparable performance to GPT-3.5-turbo in tool usage evaluations. Chen et al. (2023a) introduce FireAct, which can fine-tune Llama-2-13B to outperform prompting GPT-3.5-turbo on HotpotQA (Yang et al., 2018). Gorilla (Patil et al., 2023), fine-tuned from Llama-7B, outperforms GPT-4 on writing API calls. More recently, OpenChat-3.5(Wang et al., 2023c), an LLM pretrained with mixed quality data, has demonstrated to outperfom GPT-3.5-turbo on several agentic tasks, according to Christianos et al. (2023).
3.3 Problem Solving and Logical Reasoning Capabilities
@@
Problem solving and logical reasoning serve as fundamental capability of high-level ability and skill, such as programming, theorem proving, as well as arithmetic reasoning. To this end, in this section, we will cover the following benchmarks:
@@
  • TheoremQA (Wenhu et al., 2023) is a theorem-driven QA dataset designed to evaluate AI models’ capabilities to apply theorems to solve challenging science problems. It was curated by domain experts containing 800 high-quality questions covering 350 theorems from Math, Physics, EE&CS, and Finance.
  • HumanEval (Chen et al., 2021a) is a set of 164 hand written programming problems. Each problem includes a function signature, docstring, body, and several unit tests, with an average of 7.7 tests per problem.
@@
  • APPs (Hendrycks et al., 2021a) is a benchmark for code generation measuring the ability of models to take an arbitrary natural language problem specification and generate satisfactory Python code. The benchmark includes 10,000 problems, which range from having simple one line solutions to being substantial algorithmic challenges.
Models GSM8K HumanEval
WizardMath-7B 54.9 -
WizardMath-13B 63.9 -
WizardMath-70B 81.6 -
WizardCoder-15B - 57.3
[…]
Lemur-70B-chat 66.3 61.0
[…]
Phi-1.5-1.3B - 41.4
GPT-3.5-turbo 57.1 48.1
GPT-4 92.0 67.0
Table 3: Model Performance on two benchmarks requiring logical reasoning, GSM8K (math) and HumanEval (coding).
@@
Different from conventional knowledge distillation based instruction tuning, Luo et al. (2023c, a) employed Evol-Instruct (Xu et al., 2023a) to construct a task-specific high quality instruction tuning dataset, where the seed instructions from Alpaca (Taori et al., 2023) have evolved to the ones either extended in knowledge boundary or the depth of task complexity. After obtaining the expanded instruction pool, the new instruction tuning dataset is generated by collecting responses from another LLM, e.g., GPT-3.5-turbo. Besides, Luo et al. (2023a) also incorporate PPO (Schulman et al., 2017) algorithm to further improve the quality of both generated instruction and answer. Specifically, they first train an instruction reward model (IRM) to rank the quality of generated evolved instructions, as well as a process-supervised reward model (PRM) to provide process supervision to intermediate reasoning process. The quality supervision is generated from a fine-tuned Llama model while the process supervision come from GPT-3.5-turbo. Finally, benefiting from the evolved depth and width of queries, the fine-tuned student model achieves even better performance than GPT-3.5-turbo. For example, WizardCoder (Luo et al., 2023c) outperforms GPT-3.5-turbo on HumanEval with 19.1% absolute improvements. WizardMath (Luo et al., 2023a) has also obtained 42.9% absolute improvements on GSM8K compared with GPT-3.5-turbo.
@@
Lemur (Xu et al., 2023d) has verified a better mixture of natural language data and code and induces the LLMs with stronger abilities on function calling, automatic programming, and agent capabilities. Specifically, Lemur-70B-chat achieves significant improvements over GPT-3.5-turbo on both HumanEval and GSM8K without task-specific fine-tuning. Phi-1 and Phi-1.5 (Gunasekar et al., 2023; Li et al., 2023f) take a different road by using the textbook as the main corpus for pre-training, which makes the strong abilities observable on much smaller language models. Furthermore, OpenChat-3.5 outperforms GPT-3.5-turbo on both GSM8K and HumanEval witha significant margin, with only 7B parameters.
@@ 3.4 Modelling Long-context Capabilities
  • SCROLLS (Shaham et al., 2022) is a popular evaluation benchmark made of 7 datasets with naturally long input. The tasks cover summarization (GovReport (Huang et al., 2021), SummScreen (Chen et al., 2021b), QMSum (Zhong et al., 2021)), question-answering (Qasper (Dasigi et al., 2021), NarrativeQA (Kociský et al., 2018), QuALITY (Pang et al., 2021)) and natural language inference (ContractNLI (Koreeda & Manning, 2021)).
  • ZeroSCROLLS (Shaham et al., 2023) builds on SCROLLS (discarding ContractNLI, reusing the 6 other datasets, and adding SQuALITY (Wang et al., 2022a), MuSiQue (Trivedi et al., 2022), SpaceDigest and BookSumSort) and only considers the zero-shot setting, evaluating LLMs out-of-the-shelf.
  • LongBench (Bai et al., 2023b) sets a bilingual English/Chinese long-context benchmark of 21 datasets across 6 tasks: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks (passage retrieval), and code completion.
@@
Model Summarization QA Agg
GR $S$ QM SQAL QPER NAQA QAL MQE SD BSS
Llama-2-70B-chat + retrieval - - 18.3 - 31.3 24.5 69.6 26.7 - -
Llama-2-long-70B-chat 26.0 15.0 20.0 20.9 52.0 31.7 82.6 27.3 55.5 46.2
GPT-3.5-turbo 21.3 16.1 15.6 20.4 49.3 25.1 66.6 27.1 49.1 49.8
GPT-3.5-turbo-16k 24.3 16.2 17.4 21.4 50.0 29.5 72.0 27.0 54.1 54.6
GPT-4 26.3 17.3 18.5 22.6 50.7 27.6 89.2 41.1 62.8 60.5
Table 4: Model performance on ZeroSCROLLS. GR is short for GovReport, $S$ is SummScreen, QM is QMSum, SQAL is SQuALITY, QPER is Qasper, NAQA is NarrativeQA, QAL is QuALITY, MQE is MuSiQue, SD is SpaceDigest, and BSS is BookSumSort.
Performance of LLMs
On LongBench, L-Eval, BAMBOO and M4LE benchmarks, GPT-3.5-turbo or its 16k version largely outperform all open-source LLMs, such as Llama-2 (Touvron et al., 2023b), LongChat (Li et al., 2023a), or Vicuna (Chiang et al., 2023) ; showing that it is not trivial to drive up the performance of open-source LLM on long-input tasks.
Llama-2-long (Xiong et al., 2023) continues pre-training of Llama-2 with 400B tokens using a 16k context window (up from 4k in Llama-2). The resulting Llama-2-long-chat-70B outperforms GPT-3.5-turbo-16k by 37.7 to 36.7 on ZeroSCROLLS. Approaches to tackle long-context tasks include context window extension with positional interpolation (Chen et al., 2023e), which involves another (short) round of fine-tuning with longer context window; and retrieval augmentation (Lewis et al., 2020), which necessitates access to a retriever to find relevant information. Xu et al. (2023b) combine both these seemingly opposite techniques, pushing a Llama-2-70B above GPT-3.5-turbo-16k on average over 7 long-context tasks (including 4 datasets from ZeroSCROLLS). We refer to Table 4 for a synthesis of results on ZeroSCROLLS.
@@ 3.5.1 Query-focused Summarization
Query-focused or aspect-based summarization requires to generate summaries with regard to a fine-grained question or an aspect category. Query-focused datasets include AQuaMuse (Kulkarni et al., 2020), QMSum (Zhong et al., 2021) and SQuALITY (Wang et al., 2022a), while Aspect-based datasets include CovidET (Zhang et al., 2023a), NEWTS (Bahrainian et al., 2022), WikiAsp (Hayashi et al., 2021), etc.
Model Performance
Yang et al. (2023e) finds that standard fine-tuning on training data is still better in performance compared to ChatGPT, with an average of 2 points ROUGE-1 improvement over CovidET, NEWTS, QMSum and SQuALITY.
@@ 3.5.2 Open-ended QA
Model Performance
InstructRetro (Wang et al., 2023a) shows large improvement over GPT-3 on NQ, TriviaQA, SQuAD 2.0 and DROP, while having 7-10 percent improvement compared to a proprietary GPT-instruct model of similar size, over a range of short-form and long-form open-ended QA datasets. Initialized from a pre-trained GPT model, InstructRetro continues pre-training with retrieval and then undergoes instruction tuning. [^7] InstructRetro is not yet open-sourced.
@@ 3.5.3 Medical
For mental health, IMHI (Yang et al., 2023d) benchmark is constructed using 10 existing mental health analysis datasets, including mental health detection: DR (Pirina & Çöltekin, 2018), CLP (Coppersmith et al., 2015), Dreaddit (Turcan & McKeown, 2019), loneliness, SWMH and T-SID x (Ji et al., 2022); mental health cause detection: SAD (Mauriello et al., 2021), CAMS (Garg et al., 2022); mental risk factors detection: MultiWD (SATHVIK & Garg, 2023), IRF (Garg et al., 2023).
@@
Performance of LLMs
For mental health, MentalLlama-chat-13B (Yang et al., 2023d) finetunes a Llama-chat-13B model on IMHI training set. MentalLlama-chat-13B model with zero-shot prompting outperforms ChatGPT with few-shot prompting or with zero-shot prompting for 9 out of 10 tasks in IMHI. Liu et al. (2023b) proposes to fine-tune a Llama checkpoint to generate impression text given radiology report findings. The resulting Radiology-Llama-2 model outperforms ChatGPT and GPT-4 by a large margin on both MIMIC-CXR and OpenI datasets.
@@ 3.5.4 Generating Structured Responses
Rotowire (Wiseman et al., 2017), Struc-Bench-Latex and Struc-Bench-HTML (Tang et al., 2023b) are three datasets aiming to generate formatted tables based on text context. Rotowire contains NBA game summaries with corresponding score tables in plain text. Struc-Bench-Latex contains NBA game summaries and tables in Latex format, while Struc-Bench-HTML includes NBA game summaries with tables in HTML format.
Performance of LLMs
@@ 3.5.5 Generating Critiques
Shepherd (Wang et al., 2023e) is a 7B model initialized from Llama-7B and trained on community collected critique data and 1,317 examples of high quality human annotated data. Shepherd generates critiques on a range of diverse NLP datasets: AlpacaFarm, FairEval, CosmosQA (Huang et al., 2019), OBQA (Mihaylov et al., 2018a), PIQA (Bisk et al., 2020), TruthfulQA and CritiqueEval. With GPT-4 as an evaluator, Shepherd wins or equals ChatGPT over 60% of the time. With human evaluators, Shepherd is almost on-par with ChatGPT.
@@ 3.6 Towards Trust-worthy AI
To ensure LLMs can be trusted by humans in real-world applications, an important consideration is their reliability. For example, concerns on hallucination (Ye & Durrett, 2022; Zhao et al., 2023a) and safety (Zhiheng et al., 2023b) could deteriorate user trust in LLMs and lead to risks in high-impact applications.
@@ 3.6.1 Hallucination
  • HaluEval (Li et al., 2023b) is a large dataset of generated and human-annotated hallucinated samples. It spans three tasks: question answering, knowledge-grounded dialogue, and text summarization.
@@
  • SummEdits (Laban et al., 2023) is a 10-domain benchmark that measures LLMs’ abilities to detect factual inconsistencies.
@@
There already exist a few surveys (Zhang et al., 2023b; Rawte et al., 2023) that investigate different proposed methodologies to reduce on hallucination in detail. Specifically, the methods that surpass the current GPT-3.5-turbo performance can be either incorporated during fine-tuning or only at inference time. Selected model performances on various benchmarks are shown in Table 5.
@@
During inference, existing techniques apply specific decoding strategies, external knowledge augmentation, and multi-agent dialogue. For decoding, Dhuliawala et al. (2023) introduce Chain-of-Verification (CoVe), where the LLM drafts verification questions and self-verify the responses, which leads to a substantial improvement on FactScore over GPT-3.5-turbo.
For external knowledge augmentation, various frameworks incorporate different searching and prompting techniques to improve GPT-3.5-turbo performance Li et al. (2023d) design Chain-of-Knowledge (CoK), which retrieves evidences from heterogeneous knowledge sources and synthesizes it before answering. Peng et al. (2023) propose LLM-AUGMENTER, which augments LLMs with a set of plug-and-play modules and iteratively revises LLM prompts to improve model responses using feedback generated by utility functions. Knowledge Solver (KSL) (Feng et al., 2023) tries to teach LLMs to search for essential knowledge from external knowledge bases by harnessing their own strong generalizability. CRITIC (Gou et al., 2023) allows LLM to validate and progressively amend their own outputs in a manner similar to human interaction with tools. Luo et al. (2023b) introduce a Parametric Knowledge Guiding (PKG) framework, which equips LLMs with a knowledge-guiding module to access relevant knowledge without altering the LLMs’ parameters. These inference techniques then improve answer accuracy compared to the naive prompting strategy using GPT-3.5-turbo. Currently, GPT-3.5-turbo has also incorporated a retrieval plugin (OpenAI, 2023a) that accesses external knowledge to reduce hallucinations.
For multi-agent dialogue, Cohen et al. (2023) facilitate a multi-turn interaction between the Examinee LLM that generated the claim and another Examiner LLM which introduces questions to discover inconsistencies. Through the cross-examination process, performance on various QA tasks is improved. Du et al. (2023) ask multiple language model instances to propose and debate their individual responses and reasoning processes over multiple rounds to arrive at a common final answer, which improves on multiple benchmarks.